TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-20 23:00:00
subject: 3\13 Pt-1 ISS On Orbit Status 13-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\13 ISS On-Orbit Status 13-03-2003
Part 1 of 2

ISS On-Orbit Status 3/13/03

All ISS systems continue to function nominally, except as noted 
previously or below.

Last night's reboost by Progress was nominal. As the first of two 
Progress dynamic tests, the firing of the DPO manifold #1 (four 
thrusters) began at 5:58pm EST, for a duration of 10 minutes. It 
produced a delta-V of 1.31 m/sec (predicted: 1.39), resulting in an 
altitude increase of 2.34 km/1.46 s.mi. (predicted: 2.48 km/1.55 
s.mi.). [Being an apogee burn, it raised perigee (by 4 km) and left 
apogee essentially unchanged; this lowered the eccentricity 
(ellipticity) of the orbit, i.e., made it more circular. The test of 
the second 4-thruster manifold is scheduled for tonight at 6:37pm EST 
(delta-V: ~0.31 m/sec; delta-altitude: ~0.54 km). The cyclogram for 
its automated command sequence has already been uplinked, and flight 
attitude handover from U.S. to Russian motion control system (MCS) is 
set for 2:55pm.  The thruster tests are conducted in Mode 1, i.e., 
using propellants from the Progress M's resupply tanks (SD).]

After crew wake-up (1:00am) and morning preps, FE-1 Nikolai Budarin
performed his daily routine inspection of the Russian BIO-5
Rasteniya-2/Lada-2 plant growth experiment. [Today's activity included
execution of programmed cyclogram steps of watering the substrate in 
the growth tray, setting cultivation mode, and storing data, followed 
by photoimage recording. The data were then downlinked and discussed 
in a tag-up with a ground specialist.]

Afterwards, Budarin moved to the FGB for another round of the periodic
cleaning of protective mesh screens of Zarya's ventilation fans, a 
one-hour preventive maintenance job.

Next, Nikolai worked several hours in the Service Module (SM), 
inspecting and photo-documenting its quartz-glass window #9 in the 
frustum section between compartments (facing toward Earth and used for 
science). [Objective of the inspection, using digital still camera and 
voice recorder, is to assess the pane surfaces for any changes (new 
cavities, scratches, new or expanded old stains or discolorations 
affecting transparency properties) since the last inspection, 
performed by Nikolai on 1/13/03. The new assessment will be compared 
to the earlier observations. Budarin had extra time reserved to study 
and rehearse uplinked procedures for measuring new visible defects on 
the window, where 16 have so far been identified (millimeter range). 
Defects are measured with the parallax method which uses 
eyeball-sighting with a ruler and a right isosceles triangle to 
determine the formations' size and position with respect to the 
window's internal surface (parallax being the apparent change in an 
object's position resulting from changing the observer's position).]

CDR Ken Bowersox and FE-2/SO Don Pettit conducted the third Robotics
operations since last week, after connecting the DCP (display & 
control panel) cable for both the Lab and Cupola RWS (robotics 
workstation), today once more surveying the S1 truss segment. [Last 
week's S1 RBVM (radiator beam valve module) survey showed two 
additional QD (quick disconnect) thermal protection "booties" as 
possibly being open. Bowersox and Pettit today "walked" the SSRMS 
(space station remote manipulator system) to the MBS PDGF1 (mobile 
base system/power data grapple fixture #1) and maneuvered the arm to 
two viewing positions to inspect these booties (one was found to be 
OK, the other remains indeterminate), followed by a general survey of 
the RBVMs and two checkouts of the Tip LEE (latching end effector). At 
the end, the SSRMS was back at the Lab PDGF, positioned for viewing 
the upcoming stage EVA #2.]

Pettit completed another installment of the continuing MSG 
(Microgravity Science Glovebox) troubleshooting, which he has been 
supporting by repeated rotations of the MSG rack and reconnecting 
cables at its electronics box (E-Box). [Once again, early analysis 
showed no off-nominal current draw on the ESEM3 (exchangeable standard 
electronic module 3) after yesterday's re-installation of the P112 and 
P113 connectors. Today, Pettit re-installed only one plug, P116, which 
powers an E-Box daughter board. This leaves only one last connector to 
re-install (P117). The putative fault has yet to show itself. The 
final cable re-connection will be made tomorrow, and while the rack is 
rotated with the back panel off, Pettit is to change out the failed 
left MSG fan. The procedure is simple and short (approx. 15 min.), and 
the rack will be powered off during this time.]

Bowersox and Pettit recorded their weekly HMS FFQ (health maintenance
systems/food frequency questionnaire), a special software log on the 
medical equipment computer to track nutritional intake.

Budarin conducted the weekly inspection of the BRPK air/condensate 
separator of the SRVK water processing system. The BRPK-1 is scheduled 
for replacement with a new unit tomorrow. [The separator contains 
porous Cermet hydrophilic (water attracting) tubes through which the 
gas-liquid mixture from the heat exchanger moves. They separate the 
air from the condensate, but when the separator exceeds its service 
life or is malfunctioning, incomplete separation of the atmospheric 
condensate occurs, and the water then collects under the "sheet" of 
porous fluoroplastic. This is the main focus of the regular BRPK 
inspection and their frequent replacement.]

Nikolai also completed the daily routine maintenance of the SM's SOZh 
life support system, including ASU toilet facilities. Later. he 
prepared the daily "delta" file of the IMS (inventory management 
system) database.

The GNC R3 software patch #2 was successfully loaded into both the 
primary and backup GNC (guidance, navigation & control) MDMs last 
night. Preliminary data indicate that the software is working 
properly. [Attitude control was resumed by USOS (U.S. segment) at 
6:41pm EST after the reboost maneuver. The current navigation 
configuration has GPS1 (global positioning system #1) providing state 
vector and attitude data to the U.S. MCS, while Russian segment (RS) 
data is prime for attitude rate. If performance with the new patch 
continues to be nominal, ISS control will soon return to a 
configuration that uses rate data provided by USOS RGA (rate gyro 
assembly).]

All crewmembers performed their daily physical exercise program on RED
(resistive exercise device), TVIS (treadmill with vibration isolation 
and stabilization), CEVIS (cycle ergometer with vibration isolation) 
and VELO with load trainer. [MCC-H specialists are looking at a 
suggestion by the CDR to use a spare HDD (hard disk drive) as a 
possible replacement for the failed hard drive in the CEVIS control 
panel.]

A procedure has been uplinked for installing new alkaline batteries in 
the CDM (carbon dioxide monitor) and CSA-CP (compound specific
analyzer-combustion products) battery packs. [These batteries can be 
used with CDM, but CSA-CP requires specialcertification for alkaline 
batteries (which should be completed by the end-March) and for now is 
using only lithium batteries.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.